OpenStudy (anonymous):

Solve: 4x − 2(x + 1) = 3x + 10 The answer is 12 right?

OpenStudy (anonymous):

Distrubutive Law

OpenStudy (anonymous):

4x - 2x - 2 = 3x + 10 2x - 3x = 12 -x = 12 x = -12

OpenStudy (anonymous):

4x-2x-2=3x+10 2x-2=3x+10 2x-3x=10+2 -x = 12 x= -12

OpenStudy (anonymous):

I thought it was 4x-2x-2= 3x+10 2x-2=3x+10 2x=3x+12 x=12

OpenStudy (anonymous):

cuz your subtracting 4x-2x so how would you 3x?

OpenStudy (anonymous):

yeah but the method is the main thing here so without the distributive law u won't be able to get the right answer

OpenStudy (anonymous):

Yea i know I distributed -2(x+1) thats how i got the 4x-2x

OpenStudy (anonymous):

I don't know how you got the 3x there on the left side of the equation

OpenStudy (anonymous):

@abayomi12